Документ взят из кэша поисковой машины. Адрес оригинального документа : http://www.stsci.edu/smov/notes/0324smov.html
Дата изменения: Tue Mar 25 16:26:53 1997
Дата индексирования: Thu Feb 27 23:55:32 2014
Кодировка:

Поисковые слова: http www.badastronomy.com bad tv foxapollo.html
SMOV Status & Planning Meeting Minutes 03/24/97

SMOV Status & Planning Meeting Minutes

Meeting Date: 03/24/97



NOTE:  THE NEXT MEETING IS WEDNESDAY, MARCH 26, 8:30 AM.


STATUS REPORTS

1.  Observatory Status - All subsystems nominal

	PCS - Two HSTARS were generated since the last meeting:
		1.  HSTAR 5990 - fine lock acquisition problems
		    at 801/2350z (WFPC2 Cycle 6 Observation)
		    Still under study, but most likely a problem
 		    with the star itself.

		2.  HSTAR 5991 - Acquisition failed due to
		    attitude error at 082/2045 (WFPC2 Cycle 6 ob-
		    servation)  A slew roll update preceeded the
		    acquisition failure.  Still being
 		    investigated.

		FGS 1R (7026) The last FF3 adjustment
		    in the y-coordinate resulted in an unexpected
		    but small decrease in x-axis S-Curve levels.
	            The recommendation is to leave the FF3 in 
	            its current position.   The FGS-1R "9 Points
		    of Light" is currently executing.


2.  WFPC2 Status - Proceeding nominally with Cycle 6 science and
	remaining SMOV monitors.  In response to a request for
	UV monitoring updates, it was said that an update on
	UV throughput would be provided at the next TIPS
	meeing (Thurday).   It was also reported that the two
	UV measurements taken since the SM desorption move show
	results that are consistent with the SM at ~ 0.9 mm, as
	expected.

3.  NICMOS Status

	- Engineering Status - The Detector temperature is
	~60.7 degK.  The VCS temperature is now ~98.8 degK
	Detector warming rates remain small and VCS cooling
 	rates also imply an approach to equilibrium.  There
	remains however a small thermal gradient across the
	N2_ice/coldwell interface.

	On Friday and Saturday, the 7054 NICMOS Thermal Check
	on COSTAR Arm Deployment executed nominally.  The 
	filter wheels rotated when commanded, with no retries.
	Following the test, ops request to transition to SAA
	Operate Mode was executed, and resulted in an HSTAR
	for a TPG1 reset failure.  (This is a known problem 
	that has been reported previously at this meeting.)  
	There were no problems after several deep SAA passages.

	Upon completion of 7054, the filter wheels were
	repositioned, via realtime commanding, to their
	scientifically optimum configuration (FW1=F110M,
	FW2=FW3=F160W).  There were no retries associated with
	these realtime operations.

	- Science Status:
	Proposal 7039 - NICMOS/FGS Alignment - completed Friday
	evening and analysis has resulted in an aperture update
	of 3-4 arcseconds.  This update also has a component
	for PAM best focus.  The update has been passed on to
	the OSG group to prepare for inclusion in the SMS090.

	Proposal 7043 - (Revised) Focus Monitor has been produced
	and will be reviewed internally today by the team,
 	followed by a PIT meeting with commandos and schedulers.

	Proposal 7054 - NICMOS Thermal Check on COSTAR Deploy.  As
	reported above the filter wheels behaved nominally.  Pre-
	liminary analysis indicates a significantly lower-than-
	expected OTA/Secondary IR emissivity (<~ 20% of
 	expectations), which is good news for NICMOS
 	long-wavelength science.  (A caution that these results
	are still preliminary.)

	
4.  STIS Status

	- Engineering Status - Nominal.  
	
	- Science Status - The CCD UV Monitor (7063) that
 	occurred last Thursday evening again shows no decrease in
 	UV throughput compared to the previous measurement.

	The ERO 7124 has been reworked, following guide star
	availability problems, the new approach stepping the
	slit across the target.   This is still planned for
	inclusion in the SMS104, although people pointed out
	that time is getting tight during that week because 
	of the 12-hour hold needed for the FSW uplink for
	STIS SAA operations management and because of the
	schedule 12-hr annealing proposal (7107).


5.  FGS status - See PCS report above.


6.  FOC/COSTAR status- no new activity.   The FOC COSTAR mirror
	arms were successfully deployment on Friday.  The x-tilt
	position was as expected.  The y-tilt was a couple of 
	counts away from the pre-SM2 postion.  The arm heaters
	are on, and were on during the post-deploy portion of
	NICMOS 7054.




PLANNING REPORTS

1.  Status of Upcoming SMSs

   	- SMS083 - Currently executing.

 		 - This SMS implements the SSR for engineering
 		   data record.  This SMS also has been modified
 		   to include the Target of Opportunity
 		   identified lst week.

	- SMS090 - Initial Mode products are available on AP7
		   for review.
		   This calendar has being rebuilt to include,
		   per Project request, additional NICMOS focus
		   monitoring activities.  (See NICMOS report
		   above.)  The focus monitor for this SMS occurs
		   on day 095.

		   The NICMOS SIAF update needed for this SMS
		   means that Final Mode products will be out
		   not before Wednesday.

	  	   A "10.2" commanding update is also needed for
		   this SMS.

		   The NICMOS ERO 7111 (OMC1) has been removed
		   from this calendar, revised to use camera 2
		   instead of camera 3, and is slated for
		   SMS097.

	- SMS097 - Being built.  This calendar also will include
		   additional focus monitoring activities.

	- SMS104 - Will require a "10.2 commanding update for
		   NICMOS SAA avoidance and for STIS CCD 
		   annealing process.  This update is on
 		   schedule.  The annealing proposal (7107) can
 		   go in this SMS.

	- SMS111 - The current plan is to include the MAMA turn-on
		   (7105 revised) in this SMS.


2.  Replans

	STIS

	- MAMA HV Turn-on (7105) is delayed (from 21 March)
	  indefinitely.  (Current plan is for SMS111)  Affects on
	  downstream MAMA scheduling will be assessed.

	NICMOS

	- None yet identified (beyond the increased focus moni-
	  toring.


OPERATIONS STATUS

1.  OPUS - Some NICMOS 7054 observations for proposal went to
 	 trouble in CALNICA because of mismatch in filter
	 specifications.  The raw data were transferred via ftp to
 	 the users.  Meanwhile, OPUS is investigating.
  

2.  Archive - Nominal.



biagetti@stsci.edu

Go Back Go back to the Meeting Minutes Page